OpenStudy (anonymous):

If someone can please help me (Don't give me the answer) If 1.6 moles of an ideal gas are at a temperature of 25°C and a pressure of 0.91 atm Hg, what volume would the gas occupy, in liters?

OpenStudy (anonymous):

@Jonask @robtobey

OpenStudy (anonymous):

@Jonask @AnElephant @azura @AravindG @dan815 @DLS @dumbsearch2 @divagirl421 @e.cociuba @eragon4 @FutureMathProfessor @Falco276 @genius12 @GoldPhenoix @ivancsc1996 @kelliegirl33 @kymber @Luigi0210 @luckythebest @merchandize @melody16 @marylou004 @mariahk19 @Nurali @NoelGreco @NANAGX14 @oldrin.bataku @robtobey @skullpatrol @stabar @SI123 @skibum22 @tcarroll010 @timo86m @thomaster @touseii45 @tonipoo610 @uri @XAY_____ @yahya90 @zacharyf @zoehaddad

OpenStudy (anonymous):

Apply the formula PV=nRT i hope

OpenStudy (ivancsc1996):

Exactly. Can you solve PV=nRT for V?

OpenStudy (anonymous):

; -; Uh... V= Pn/RT?

OpenStudy (anonymous):

V= nRT/P

OpenStudy (ivancsc1996):

No, you only need to divide by V so P=nRT/P. Then jjust plug in the numbers. remember you need the temperature in kelvin

OpenStudy (anonymous):

If 1.6 moles of an ideal gas are at a temperature of 25°C and a pressure of 0.91 atm Hg, what volume would the gas occupy, in liters? Okay so P = 0.91 T= 25°C n= ??? R= ???

OpenStudy (ivancsc1996):

n is the number of moles. R is the gas constant, you can look for it on google. remember T is T=(25+273.15)k

OpenStudy (anonymous):

Okay so... How do I fined the gas content or is it the same as the moles?

OpenStudy (anonymous):

\[0.0821 L atm mol ^{-1} /K ^{-1}\]

OpenStudy (anonymous):

e__e okay so the answer will be 43L?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you! c: <3